OpenStudy (anonymous):

can somebody explain me implicit differentiation with this problem xy^3 + xy = 4

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

\[\begin{align*}\frac{d}{dx}\left[xy^3+xy\right]&=\frac{d}{dx}4\\ \frac{d}{dx}[xy^3]+\frac{d}{dx}[xy]&=0\\ \frac{d}{dx}[x]y^3+x\frac{d}{dx}[y^3]+\frac{d}{dx}[x]y+x\frac{d}{dx}[y]&=0\\ y^3+x\left(3y^2\frac{dy}{dx}\right)+y+x\frac{dy}{dx}&=0 \end{align*}\]

OpenStudy (anonymous):

Solve for \(\dfrac{dy}{dx}\)

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@SithsAndGiggles could you explain it to me in words, please and thank you

OpenStudy (imstuck):

i got the answer but it is so difficult to explain implicit diff. Hopefully @zzr0ck3r can do a fair job of it!

OpenStudy (anonymous):

the answer is DAK DAK DAK DAK DAK DAK DAK

OpenStudy (zzr0ck3r):

line 1: take derivative of both sides line 2:use the fact that differentiation is a linear operator so you can "split" things apart line3:use the product rule for differentiating line 4:take all the derivatives note: since y depends on x, when we take the derivative of 2y you get 2 y' and the derivative of y^2 = 2yy' because of the chain rule

OpenStudy (dan815):

chain rule, since y=f(x) so if you have g(y) then dg/dx= g'(y)*dy/dx

OpenStudy (zzr0ck3r):

and holy god is @SithsAndGiggles do a good job of making it look pretty

OpenStudy (anonymous):

@zzr0ck3r thanks!

OpenStudy (dan815):

agreed lol